# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v2

@@ 7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 2, 7, and 10). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 21).
@@ 7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 8). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 6 and 10). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 4), supporting accessibility and inclusivity concerns (Implication 11), and enabling more personalized and flexible user experiences (Implication 7). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 13).
@@ 7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis approach provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality.
@@ 7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 4). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 4), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
@@ 8. Threats to Validity
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
@@ 9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,000 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
@@
Data Availability. We provide a replication package, including manually labelled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.

### v3

@@ 7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 1, 3, and 5). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 5).
@@ 7.2. Implications for Platform Owners
Beyond improving review quality, our findings also reveal structural limitations in how app stores currently support developer–user interaction. As shown in our developer response analysis (Implication 6), review threads in Gen-AI apps function primarily as an initial point of contact rather than a space for resolving AI-specific failures. Many issues require diagnostic information that cannot be shared publicly or involve model- and policy-level changes that cannot be deployed quickly. Platform owners can play an important role by offering richer, privacy-preserving support channels that complement review sections and allow developers to investigate Gen-AI issues more effectively. Strengthening these communication pathways would improve user experience and reduce the friction developers face when addressing complex AI behaviours.
@@ 7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 4). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 3 and 5). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 2), supporting accessibility and inclusivity concerns (Implication 5), and enabling more personalized and flexible user experiences (Implication 3). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 7).
@@ 7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis framework provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality. Beyond Gen-AI apps, the same framework and analytical tools can be applied to user reviews in other domains, enabling engineers to systematically identify evolving requirements and usability concerns across a wide range of software systems.
@@ 7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 2). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 2), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
@@ 8. Threats to Validity
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. This single-platform scope may limit the generalizability of results to other ecosystems with different user bases or review practices. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
@@ 9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,066 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
@@
Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at omitted.